OpenStudy (anonymous):

What is an equation of the line,in point -slope form, that passes through the given point and ha the given slope Points(8,-8) Slope:3 Answers are Y-8=3(x+8) Y+8=3(x+8) Y-8=3(x-8) Y+8=3(x-8) Thanks

OpenStudy (vshiroky):

y-y1=3(x-x1)

OpenStudy (vshiroky):

y-(-8)=3(x-8) negative and negative= pos